ou may need to use the appropriate appendix table to answer this question.

Suppose that the mean daily viewing time of television is 8.35 hours. Use a normal probability distribution with a standard deviation of 2.5 hours to answer the following questions about daily television viewing per household

a.How many hours of television viewing must a household have in order to be in the top 2% of all television viewing households? (Round your answer to two decimal places.)

b.What is the probability that a household views television more than 3 hours a day? (Round your answer to four decimal places.)
Answer #1

Given that,

(a)

P(Z > 2.33) = 0.02

z = 2.33

Using z-score formula,

x = z * +

x = 2.33 * 2.5 + 8.35 = 14.17

Answer = 14.17 hours

(b)

P(x > 3) = 1 - P(x < 3)

= 1 - P[(x - ) / < (3 - 8.35) / 2.5]

= 1 - P(z < -2.14)

= 1 - 0.0162

= 0.9838

Probability = 0.9838